OpenStudy (zappy620):

Find the derivative of y=cotx/(1+cotx)

OpenStudy (zappy620):

\[y= \frac{cotx}{1+cotx}\]

OpenStudy (zappy620):

Sure :D

OpenStudy (zappy620):

Im just trying to show how far I gotten with the problem

OpenStudy (zappy620):

\[\frac{(1+cotx)(-\csc^2x)-(cotx)(-\csc^2x)}{(1+\cot^2)}\]

OpenStudy (steve816):

I see one error, on the denominator, the squared should be in the outside of the parenthesis. (1 + cotx)^2

OpenStudy (zappy620):

oh yea you're right my mistake

OpenStudy (steve816):

Otherwise, you did the quotient rule right and is just a matter of simplifying!

OpenStudy (zappy620):

\[\frac{(1+cotx)(-\csc^2x)-(cotx)(-\csc^2x)}{(1+\cot)^2}\]

OpenStudy (solomonzelman):

\(\color{black}{ \displaystyle y=\frac{\cot x}{1+\cot x}=\frac{\sin x}{\sin x+\cos x}}\). (I would differentiate then.)

OpenStudy (zappy620):

I'm just i guess intimidated on how to simplify it honestly. I don't know what cot*csc^2 is

OpenStudy (zappy620):

Did you just use the ratios to simply the given equation?

OpenStudy (zappy620):

\[\frac{(1+cotx)(-\csc^2x)-(cotx)(-\csc^2x)}{(1+\cot)^2}\] I'm still stick on figuring out how to simplify this :( And I don't know how solomon got to the alternate form of the equation.

zepdrix (zepdrix):

\[\large\rm \frac{(1+cot x)\color{orangered}{(-\csc^2x)}-(cot x)\color{orangered}{(-\csc^2x)}}{(1+\cot x)^2}\]We could try factoring this orange thing out of each term, ya?

zepdrix (zepdrix):

Multiplying everything together seems like a messy alternative :P

OpenStudy (zappy620):

:O

OpenStudy (zappy620):

I didn't even see that omg thank you!

OpenStudy (zappy620):

I'll continue the problem on paper then type the work up here.

zepdrix (zepdrix):

k

OpenStudy (zappy620):

Did I factor it out right? \[\frac{(-\csc^2x)(1+cotx)-(cotx)}{(1+\cot^2)}\]

zepdrix (zepdrix):

The -csc^2x should be multiplying `everything` in the numerator, so you might want another set of brackets to make that clear,\[\large\rm \frac{(-\csc^2x)\left[(1+cotx)-(cotx)\right]}{(1+\cot^2)}\]

OpenStudy (zappy620):

Ah ok

zepdrix (zepdrix):

Oh woops we've got that denominator messed up again :p\[\large\rm \frac{(-\csc^2x)\left[(1+cotx)-(cotx)\right]}{(1+\cot x)^2}\]

OpenStudy (zappy620):

Haha yea I just copy and pasted it my bad.

OpenStudy (zappy620):

So would the cotx cancel out in the numerator?

zepdrix (zepdrix):

\[\large\rm \frac{(-\csc^2x)\left[1\right]}{(1+\cot x)^2}\]Yes.

OpenStudy (zappy620):

Alright :D So the denominator would equal 1 due to the identity right?

OpenStudy (zappy620):

Sorry internet is being kind of slow right now...

zepdrix (zepdrix):

Equals 1? Hmm I'm not sure what you mean..

OpenStudy (zappy620):

Oh no i read it wrong...

zepdrix (zepdrix):

Remember that our Pythagorean Identity looks like this,\[\large\rm 1+\cot^2x=\csc^2x\]Not like this,\[\large\rm (1+\cot x)^2\ne \csc^2x\]

zepdrix (zepdrix):

If that's maybe what you were referring to.

zepdrix (zepdrix):

From this point, I actually don't see a nice path using any cosecant or cotangent identity. I think it's time to convert to sines and cosines and continue on from there.

OpenStudy (zappy620):

Ok do you usually convert to sin and cos at the end?

zepdrix (zepdrix):

Honestly, it just depends what your teacher expects. This is a fine and dandy answer:\[\large\rm \frac{-\csc^2x(1+\cot x)-\csc^2x(\cot x)}{(1+\cot x)^2}\] This answer is also good:\[\large\rm \frac{-\csc^2x}{(1+\cot x)^2}\]Or in terms of sines and cosines will also give a nice looking answer. Just depends.

zepdrix (zepdrix):

Based on problems you've already done, you should have a good idea of what your teacher expects for a good looking answer. Hopefully :d hehe

OpenStudy (zappy620):

Ok well in the book it has the answer converted to sin and cos actually.

zepdrix (zepdrix):

Ok ok fair enough :d so we'll keep going then.

OpenStudy (zappy620):

So for converting im having trouble visualizing it. This is what I have right now \[\frac{ (1)(\sin) }{ (sinx^2+\cos +1) }\]

OpenStudy (zappy620):

Forgot the negative sign

OpenStudy (zappy620):

and forgot to square the denominator...

zepdrix (zepdrix):

Hmm I was thinking of it like this,\[\large\rm \frac{\color{orangered}{-\csc^2x}}{(1+\color{royalblue}{\cot x})^2}\quad=\quad \frac{\color{orangered}{-\frac{1}{\sin^2x}}}{\left(1+\color{royalblue}{\frac{\cos x}{\sin x}}\right)^2}\]

OpenStudy (zappy620):

Oh ok. Still the fraction ception is kinda messing me up still.

zepdrix (zepdrix):

This might seem like a weird step, you can lemme know if it's confusing... In order to get rid of these fractions upon fractions, we need to multiply the top and bottom by sin^2x.

OpenStudy (zappy620):

Oh ok to cancel it out i see.

zepdrix (zepdrix):

\[\large\rm \frac{\left(-\frac{1}{\sin^2x}\right)}{\left(1+\frac{\cos x}{\sin x}\right)^2}\left(\frac{\sin^2x}{\sin^2x}\right)\]

OpenStudy (zappy620):

So your left with -1 in the numerator. I'm not sure about the denominator since its sinx*sin^2x

zepdrix (zepdrix):

Ok bottom is a little weird.\[\large\rm \sin^2x\left(1+\frac{\cos x}{\sin x}\right)^2\]We have to drop the square on our sine in order to bring it into this other square.

zepdrix (zepdrix):

Example of this type of idea:\[\large\rm 2^2(3)^2\quad=\quad (2\cdot3)^2\]I had to drop the square on that first 2 in order to bring it into the other group. Hope that's not too confusing.

OpenStudy (zappy620):

Oh ok.

zepdrix (zepdrix):

\[\large\rm \sin^2x\left[\left(1+\frac{\cos x}{\sin x}\right)\right]^2\quad=\quad \left[\sin x\left(1+\frac{\cos x}{\sin x}\right)\right]^2\]

OpenStudy (zappy620):

I see. So the denominator would be 1+cosx right?

OpenStudy (zappy620):

Oh no (1+cos)^2

zepdrix (zepdrix):

Woops, you didn't multiply the 1 by sinx.

zepdrix (zepdrix):

Make sure you distribute the sinx to each term :)

OpenStudy (zappy620):

Oh i see my bad. It would be (sinx + cosx)^2

zepdrix (zepdrix):

Mmm good good, that looks better.

zepdrix (zepdrix):

Oh I guess we could simplify further :\ Ughhh

OpenStudy (zappy620):

Actually no. The answer we got so far is \[\frac{ 1 }{ (sinx+cosx)^2 }\]

OpenStudy (zappy620):

Right*?

OpenStudy (zappy620):

I'm really sorry i meant it like that's the answer in the book.

zepdrix (zepdrix):

\[\large\rm \frac{-1}{(sinx+cosx)^2}\]with a negative up top? Yes :) looks good. If that's where your book stopped, then that's fine. I was just going to point out that we COULD distribute the square in the bottom,\[\large\rm (\sin x+\cos x)^2\quad=\quad \sin^2x+2\sin x \cos x+\cos^2x\]And then apply your Pythagorean Identity with the sin^2x+cos^2x,\[\large\rm =1+2\sin x \cos x\]

OpenStudy (zappy620):

Oh i see lol. I didn't think about it like that actually

zepdrix (zepdrix):

Boy that was a doozy :P But we did it! Yayyyy! \c:/ Zeppy and Zappy!

OpenStudy (zappy620):

Omg thank you so much for helping me through this problem. I literally spent like prolly 3 hours now on that problem.. I honestly really appreciate you for stick with me. Thank you.

zepdrix (zepdrix):

np

zepdrix (zepdrix):

If that 'multiplying top and bottom by sin^2x` was a confusing step, you had another option. You could instead look for a common denominator in your denominator, in the square, and the apply your weird "keep change flip" rule or however you learned it, flipping the bottom fraction, rewriting it as multiplication. And the sin^2x's would have cancelled out that way. But multiplying through by LCM of denominators is a really useful tool to have in your goody bag, so I wasn't to make sure we at least touched on it.

OpenStudy (zappy620):

Ah ok. To be honest I liked the way you first showed it. I was just having a super hard time visualizing the conversion on paper.

OpenStudy (zappy620):

I think its a good time to take a break from math for a bit. My brain is kinda having a hard time functioning right now lol. But, I can't thank you enough for helping me with this problem. You truly are a kind person :D